could you answer, please asap

media%2Fb94%2Fb947400c-05d6-4bfd-9922-82

could you answer, please asap
Answer #1

a) initial height = 20 m

b) maximum height reached = 100 m

time taken to reach maximum height = 4 seconds

height travelled = 100-20=80 m

let initial velocity be u

80=u*4 - 0.5*10*42=4u-80

u=40 m/s

c)maximum height = 100 m

d)time taken to fall through 100 m = 8.5-4=4.5 seconds

final velocity = 10*4.5=45 m/s

e)total time of trip = 8.5 seconds